OpenStudy (anonymous):

What is the midpoint of (–4, 4) and (2, –4)? (–2, 0) (–6, 8) (0, 0) (–1, 0)

OpenStudy (anonymous):

(-1,0)

OpenStudy (anonymous):

\( (–1, 0) \)

OpenStudy (anonymous):

Its just x=(-4+2)/2=-1 and y=(-4=4)/2=0 And point is(-1,0)

OpenStudy (anonymous):